OpenStudy (anonymous):

Write the point-slope equation of the line through (-2, 1) with a slope of 3. A. y - 1 = 3(x + 2) B. y - 2 = 3(x + 1) C. y + 2 = 3(x - 1)

OpenStudy (anonymous):

Do you need help?

OpenStudy (anonymous):

y - 1 = 3(x + 2)

OpenStudy (anonymous):

1) Use the formula y-y1=m(x-x1) to find your equation 2) Your eqaution would look like y-1=3(x+2) 3) Distribute to get y-1=3x+6 4) Add 1 to both sides to get y=3x+7 5) In your case the answer would be (A) Hope that helps :D